-- JoergStelzer - 2022-08-10

Run number Date Length Fill System with problems Comments e-logs and jiras
427394 Tue, Jul 5   2b   first collisions of Run 3  
427405 Tue, Jul 6   2b      
427514 Tue, Jul 7   8b      
427882 Tue, Jul 10   62b      
427883 Tue, Jul 10   62b      
427884 Tue, Jul 10   62b      
427885 Tue, Jul 10   62b      
427892 Tue, Jul 11   62b      
427911 Tue, Jul 11   62b      
427914 Mon, Jul 11   62b      
427927 Tue, Jul 12   62b      
427929 Tue, Jul 12   64b, mu=0.005   very low mu run  
428071 Tue, Jul 12   326b, mu=39   12b-trains. Toroid off.  
428353 Thu, Jul 14   302b, mu=32   48b-trains  
428580 Sun, Jul 17   302b, mu=37      
428648 Tue Jul 19   590b, mu=35      
428700 Wed, Jul 20   590b, mu=1 AFP low-mu run for AFP. TOF readout off by 1 BC.  
428747 Wed, Jul 20   590b, mu=37 AFP low-mu run for AFP. TOF readout off by 1 BC.  
428759 Wed, Jul 20   590b, mu=0.05 AFP low-mu run for AFP. TOF readout off by 1 BC.  
428770 Wed, Jul 20   590b, mu=0.05   low-mu run for AFP  
428776 Wed, Jul 20   590b, mu=32      
428777 Thu, Jul 21   590b, mu=30      
428855 Thu, Jul 21   590b, mu=53 TRT TRT resync storm above mu=43. First occurance.  
429018 Thu, Jul 21   974b, mu=39      
429027 Sat, Jul 23   974b, mu=39      
429137 Mon, Jul 25   974b, mu=36      
429142 Mon, Jul 25   974b, mu=50 TRT TRT resync storm about mu=46.  
429452 Wed, Jul 27   974b, mu=36      
429469 Wed, Jul 27   733b, mu=32      
429470 Thu, Jul 28   974b, mu=37      
429603 Thu, Jul 28   974b, mu=38 HLT fix first run with fix of bug in CPS logic. Bug caused an unwanted effective additional HLT prescale for HLT chains inside a CPS group ATR-26072
429606 Thu, Jul 28   974b, mu=49      
429612 Fri, Jul 29   1214b, mu=44      
429658 Fri, Jul 29   1214b, mu=44      
429697 Sat, Jul 30   1214b, mu=43      
429716 Sun, Jul 31   1214b, mu=41      
429782 Sun, Jul 31   1538b, mu=42 HLT fix first run where HLT prescale bit information is correct ATR-26036
429929 Mon, Aug 1   1538b, mu=43      
429940 Mon, Aug 1   1538b, mu=46 HLT Testing HLT: HLT farm backpressure from ROSes at mu=45  
429993 Tue, Aug 2   1538b, mu=42      
430036 Tue, Aug 2   1538b, mu=41      
Run number Date Length Fill System with problems Comments e-logs and jiras
430178 Wed, Aug 3   1922b, mu=42      
430183 Thu, Aug 4   1922b, mu=42 PIX Pixel used 30 minutes data at end-of-fill to investigate loss of cluster efficiency  
430336 Thu, Aug 4   1922b, mu=45      
430341 Fri, Aug 5   1922b, mu=50      
430488 Fri, Aug 5   1922b, mu=44      
430490 Sat, Aug 6   1922b, mu=43      
430526 Sat, Aug 6   1922b, mu=50      
430536 Sat, Aug 6   1922b, mu=52 MDT fix first run with correct MDT online conditions and good HLT muon efficiency elog 467550, ATR-25968
430542 Sun, Aug 7   1922b, mu=50      
430580 Sun, Aug 7   1922b, mu=51      
430647 Mon, Aug 8 7m 2160b, mu=28      
430648 Tue, Aug 9 9h15m 2160b, mu=55      
430702 Tue, Aug 9 7.5h 2160b, mu=56 MUCTPI No endcap muons due to MUCTPI latency misconfiguration elog 468211, ATR-26146, elog 468388elog 468264
430896 Wed, Aug 10 52m 2160b, mu=56 Tile Low muon efficieny in eta-region 1.3. Wrong settings of TMDB, reverted for the next run https://atlasop.cern.ch/elisa/display/468545
430897 Wed, Aug 10 10h 2160b, mu=52   Farm busy at initial mu=54. Ask LHC for 52, then all fine.  
431037 Thu, Aug 11 10h 2160b, mu=52   Temporarily reached mu=56 after beta* leveling and optimization  
431178 Fri, Aug 12 4h 2400b, mu=44 RPC RPC power supply, small holes in L1 muon for first hour. TTC restart fixed it.  
431178 Sat, Aug 13 53m 2400b, mu=47      
431215 Sat, Aug 13 13h25m 2400b, mu=54      
431228 Sun, Aug 14 10m 2400b, mu=35      
431241 Sun, Aug 14 10.5h 2400b, mu=55   HLT: first run with fix for beam spot with updated ID alignment conditions (affects b-jet and perhaps tau trigger) ATR-26018
431341 Mon, Aug 15 2h15m 2400b, mu=54      
431493 Wed, Aug 17   1922b, mu=54      
431810 Fri, Aug 19 3h 2400b, mu=55 LAr fix First run with fix of LAr pileup noise normalization elog 470001, ATR-26108
431812 Fri, Aug 19   2400b, mu=55 AFP Lost 1/2 hour due to AFP LV board losing power  
431850 Sat, Aug 20 33m 2400b, mu=51      
431885 Sat, Aug 20 10h20m 2400b, mu=56 L1Calo, Pixel, HLT L1Calo 100% busy the first few LBs, due to late TTC restart. HLT farm busy, Pixel with small busy. LAR DT SWROD enabled for the first time with 100kHz.  
431894 Sun, Aug 21 4h55m 2400b, mu=52 LAR DT 15% busy from farm due to LAR SWROD not responding and not being kill at CONNECTED 470703
431906 Sun, Aug 21 10h24m 2400b, mu=52 AFP ADP lost data in 1 arm, but no way to recover (no stopless removal)  
432180 Mon, Aug 22 12h23m 2400b, mu=46 RPC 36b-trains. Lost 22 minutes due to massive RPC resync (sequential) elog 471097

Run number Date Length Fill System with problems Comments e-logs and jiras
435777 Fri, Sep 30   1154, mu=48 MDT-EC    
435816 Sat, Oct 01   2448, mu=45      
435831 Sun, Oct 02   2448, mu=40      
435854 Sun, Oct 02   2448, mu=52 LAR New OFC after Toroid ramp applied elog 476845
435931 Mon, Oct 03 9h00m 2400b, mu=51 MUCTPI MUCTPI input misconfiguration for first 4h, causing holes in the endcap muons elog 477330
435946 Tue, Oct 04   2400b, mu=52      
435981 Tue, Oct 04   12b      
436041 Wed, Oct 05   2400b, mu=55 HLT Readout limitations at high pileup, expert called late  
440485 Thu, Nov 24 4h 9b, mu=20..55   BSRT fill with extended emittance scans  
Edit | Attach | Watch | Print version | History: r25 < r24 < r23 < r22 < r21 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r25 - 2022-11-25 - JoergStelzer
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback